# Hyper-optimized Quantum Lego Contraction Schedules

### v2

@@ 6 Performance results
For most of the layouts using a universal representation (MSP and Tanner), tensor network contraction performs worse than directly calculating the WEP through brute force. The exceptions are the RSC MSP with $n=25$ and RSC Tanner with $n=25$ layouts, where the SST cost function decidedly reduces the contraction cost to below that of brute force, while the dense tensor cost function results in higher than brute force cost. For all other MSP and Tanner representations, both the SST and default cost functions perform worse than brute force, and therefore, even though SST would improve the contraction cost significantly, using brute force would be faster.
@@
Surprisingly, despite consisting of significantly more tensors, the MSP networks outperform the Tanner networks for the largest members of all three code families (rotated surface code, quantum Hamming, and bivariate bicycle codes).
@@ 7 Conclusions and outlook
The SST cost function is efficient, scales as $O(n^{3})$, and is exact, and thus can help decide whether calculating the WEPs for a given code and QL layout is better via brute force methods or using a given contraction schedule. Most of our Tanner and MSP layout examples for our 64-trial cutoff are beyond the cost of brute force. It is possible that increasing the sample size of the Cotengra hyperoptimizer would decrease the number of operations. The MSP layout, in particular, features numerous small tensors, resulting in a larger space for the hyperoptimizer to explore. Consequently, more trials could improve the quality of the contraction trees. It is unclear whether an optimal layout would be able to improve the cost below brute force. The entanglement structure of the code may render them non-viable for QL contraction. Namely, volume-law entanglement in these codes would cause beyond-brute force contraction costs [17]. There is recent evidence for the BB codes having worse-than-area-law entanglement [36], which could explain the worse-than-brute force contraction costs. The analysis and quantification of contributing factors are of great interest and will be the subject of follow-up work.
@@ Appendix D Performance Metrics for SST and Dense Cost Functions
RSC Tanner 9 $8.28\text{\times}{10}^{02}$ $7.33\text{\times}{10}^{02}$ $1.129\text{\times}{10}^{00}$
RSC Tanner 25 $2.87\text{\times}{10}^{07}$ $6.57\text{\times}{10}^{06}$ $4.374\text{\times}{10}^{00}$
RSC MSP 9 $2.21\text{\times}{10}^{03}$ $2.09\text{\times}{10}^{03}$ $1.054\text{\times}{10}^{00}$
RSC MSP 25 $5.99\text{\times}{10}^{07}$ $4.28\text{\times}{10}^{06}$ $1.399\text{\times}{10}^{01}$
@@
Concat Rep 27 $3.26\text{\times}{10}^{02}$ $3.25\text{\times}{10}^{02}$ $1.004\text{\times}{10}^{00}$
Concat Rep 81 $24\text{\times}{10}^{03} 1.1$$1\text{\times}{10}^{03} 1.1$$113\text{\times}{10}^{00}$
Holographic 25 $3.40\text{\times}{10}^{03}$ $2.70\text{\times}{10}^{03}$ $1.260\text{\times}{10}^{00}$
Holographic 95 $2.58\text{\times}{10}^{06}$ $2.02\text{\times}{10}^{06}$ $1.279\text{\times}{10}^{00}$
RSC 9 $3.36\text{\times}{10}^{02}$ $1.40\text{\times}{10}^{02}$ $2.400\text{\times}{10}^{00}$
RSC 25 $1.99\text{\times}{10}^{03}$ $8.76\text{\times}{10}^{02}$ $2.274\text{\times}{10}^{00}$
RSC 49 $1.50\text{\times}{10}^{04}$ $5.07\text{\times}{10}^{03}$ $2.961\text{\times}{10}^{00}$
RSC Tanner 9 $8.18\text{\times}{10}^{02}$ $7.43\text{\times}{10}^{02}$ $1.100\text{\times}{10}^{00}$
RSC Tanner 25 $9.44\text{\times}{10}^{06}$ $7.52\text{\times}{10}^{06}$ $1.254\text{\times}{10}^{00}$
RSC MSP 9 $54\text{\times}{10}^{03} 2.2$$28\text{\times}{10}^{03}$ $1.113\text{\times}{10}^{00}$
RSC MSP 25 $2.70\text{\times}{10}^{08}$ $6.21\text{\times}{10}^{06}$ $4.347\text{\times}{10}^{01}$

### v3

@@ 6 Performance results
For most of the layouts using a universal representation (MSP and Tanner), tensor network contraction performs worse than directly calculating the WEP through brute force. The exceptions are the RSC MSP with $n=25$ and RSC Tanner with $n=25$ layouts, where both the SST cost function and the dense tensor cost function reduce the contraction cost to below that of brute force. For all other MSP and Tanner representations, both the SST and default cost functions perform worse than brute force, and therefore, even though SST would improve the contraction cost significantly, using brute force would be faster.
@@
Surprisingly, despite consisting of significantly more tensors, the MSP networks outperform the Tanner networks in the cases of $n=15$ quantum Hamming code and the $n=18$ bivariate bicycle codes.
@@ 7 Conclusions and outlook
The SST cost function is efficient, scales as $O(n^{3})$, and is exact, and thus can help decide whether calculating the WEPs for a given code and QL layout is better via brute force methods or using a given contraction schedule. Most of our Tanner and MSP layout examples for our 64-trial cutoff are beyond the cost of brute force, except for the case of the larger, $n=25$ rotated surface code instances. It is possible that increasing the sample size of the Cotengra hyperoptimizer would decrease the cost of contraction. The MSP layout, in particular, features numerous small tensors, resulting in a larger space for the hyperoptimizer to explore. Consequently, more trials could improve the quality of the contraction trees. It is unclear whether an optimal layout would be able to improve the cost below brute force. The entanglement structure of the code may render them non-viable for QL contraction. Namely, volume-law entanglement in these codes would cause beyond-brute force contraction costs [5]. There is recent evidence for the BB codes having worse-than-area-law entanglement [38], which could explain the worse-than-brute force contraction costs. Furthermore, the area law scaling of the rotated surface code might be the very reason for the RSC contraction cost’s better-than-brute force scaling for these universal layouts. The analysis and quantification of contributing factors are of great interest and will be the subject of follow-up work.
@@ Appendix D Performance Metrics for SST and Dense Cost Functions
RSC Tanner 9 $3.19\text{\times}{10}^{02}$ $3.00\text{\times}{10}^{02}$ $1.065\text{\times}{10}^{00}$
RSC Tanner 25 $5.30\text{\times}{10}^{03}$ $3.29\text{\times}{10}^{03}$ $1.610\text{\times}{10}^{00}$
RSC MSP 9 $30\text{\times}{10}^{03} 1.1$$19\text{\times}{10}^{03}$ $1.096\text{\times}{10}^{00}$
RSC MSP 25 $8.86\text{\times}{10}^{03}$ $6.58\text{\times}{10}^{03}$ $1.347\text{\times}{10}^{00}$
@@
Concat Rep 9 $6.200\text{\times}{10}^{01}$ $6.200\text{\times}{10}^{01}$ $1.000\text{\times}{10}^{00}$
Concat Rep 27 $3.269\text{\times}{10}^{02}$ $3.257\text{\times}{10}^{02}$ $1.004\text{\times}{10}^{00}$
Concat Rep 81 $242\text{\times}{10}^{03} 1.1$$16\text{\times}{10}^{03} 1.1$$113\text{\times}{10}^{00}$
Holographic 25 $3.408\text{\times}{10}^{03}$ $2.705\text{\times}{10}^{03}$ $1.260\text{\times}{10}^{00}$
Holographic 95 $2.585\text{\times}{10}^{06}$ $2.021\text{\times}{10}^{06}$ $1.279\text{\times}{10}^{00}$
RSC 9 $3.360\text{\times}{10}^{02}$ $1.400\text{\times}{10}^{02}$ $2.400\text{\times}{10}^{00}$
RSC 25 $1.993\text{\times}{10}^{03}$ $8.762\text{\times}{10}^{02}$ $2.274\text{\times}{10}^{00}$
RSC 49 $1.503\text{\times}{10}^{04}$ $5.078\text{\times}{10}^{03}$ $2.961\text{\times}{10}^{00}$
RSC Tanner 9 $3.30\text{\times}{10}^{02}$ $3.25\text{\times}{10}^{02}$ $1.013\text{\times}{10}^{00}$
RSC Tanner 25 $3.71\text{\times}{10}^{03}$ $3.25\text{\times}{10}^{03}$ $1.140\text{\times}{10}^{00}$
RSC MSP 9 $1.38\text{\times}{10}^{03}$ $1.32\text{\times}{10}^{03}$ $1.049\text{\times}{10}^{00}$
RSC MSP 25 $1.28\text{\times}{10}^{04}$ $1.02\text{\times}{10}^{04}$ $1.258\text{\times}{10}^{00}$